# Phase Transition for Budgeted Multi-Agent Synergy

### v1

@@ 5.5 A universal upper bound on amplification exponents
Implications for synergy.
The bound $s\leq 1/2$ has two immediate consequences. First, it explains why scale-out is fragile on some tasks. If the single-agent scaling exponent satisfies $\beta\geq 1/2$, then the inequality $s>\beta$ is impossible under majority-vote hierarchies with one-bit messages. In that regime, a designer should not expect budget-exponent gains from adding agents; improvement must come from making agents stronger, from richer communication than a single bit, or from protocols that effectively reduce shared failures. Second, the bound clarifies the role of diversity. Even if $\beta<1/2$, a large $\rho$ reduces $\alpha_{\rho}$ and pushes $s$ downward, shrinking or eliminating the synergy region in Figure 8. This is precisely the regime in which practitioners observe that “more agents does not help,” and it suggests a concrete remedy: invest budget not in more of the same agents, but in mechanisms that reduce shared error modes, which in our framework means reducing the effective $\rho$.
Summary of the phase-diagram logic.
Taken together, the results above reduce topology choice to a few quantitative checks. Under context constraints, deep trees are only viable in the supercritical regime $\alpha_{\rho}>1$; exponent-level budgeted synergy in the growth range requires $s>\beta$ (equivalently $\alpha_{\rho}>b^{\beta}$); and within the majority-vote tree family we always have the universal limit $s\leq 1/2$. Even when these conditions hold, depth helps only up to saturation: once the recursion is near $\mu^{\star}$, additional budget is typically better spent improving communication fidelity (larger $\gamma(m)$) or reducing shared failures (smaller $\rho$) than on additional depth. Section 6 turns these checks into a compact design diagnostic within the model.
@@ 6 Design Diagnostics Within the Model
The results so far characterize when different organizations can help under a fixed budget: a context-limited star saturates (Section 5.3), chains compound loss (Section 5.2), and trees are viable only in the supercritical regime $\alpha_{\rho}>1$ with small-signal growth governed by an organization exponent $s$ (Section 4). Here we translate these results into design diagnostics within the model: given a budget $B$, a context window $W$, and a small set of effective environment quantities, we ask whether hierarchy is feasible, whether scale-out can outpace scale-up before saturation, and which constraint is binding when it cannot. For LLM-based agent systems, the knobs correspond to token allocations: per-leaf reasoning/compute $x$ and inter-agent message length $m$. The abstraction is deliberately coarse, so the goal is regime prediction and diagnostic guidance, not a faithful end-to-end optimizer for any particular agent stack.
@@
6.1 Closed-form compute allocation: optimal $x^{\star}$ in the growth regime
Fix a candidate tree architecture specified by a fan-in $b$ and message length $m$, with local feasibility constraint $bm\leq W$. Let $\rho$ be the calibrated shared-failure correlation and $\gamma(m)$ be the calibrated communication reliability. Define the per-layer weak-signal gain
$\alpha_{\rho}(b,m):=\gamma(m)\big(\rho+(1-\rho)f_{b}^{\prime}(0)\big),\qquad s(b,m):=\frac{\log\alpha_{\rho}(b,m)}{\log b}\quad\text{(for fixed calibrated }\rho\text{)}.$
Theorem 4 implies that deep hierarchies are only viable if $\alpha_{\rho}(b,m)>1$. Assume we are in this supercritical regime and consider budgets for which the system remains in the growth regime before saturation. Under the small-signal scaling approximation $g(x)\approx kx^{\beta}$ and the tree budget model $B\approx N(x+c_{0}(b,m))$ with $c_{0}(b,m)=2m\frac{b}{b-1}$, Section 5.4 derived the growth objective
$\mu_{\mathrm{grow}}(B;b,m,x)\;\approx\;k\,x^{\beta}\left(\frac{B}{x+c_{0}(b,m)}\right)^{s(b,m)}.$ (40)
The next fact is the inner workhorse of the design procedure: once $(b,m)$ are fixed, the optimal compute per leaf has a closed form.
Proposition 12 (Growth-regime compute allocation)
Fix $(b,m)$ with $s(b,m)>\beta$. Then the objective (40) is maximized at
$x^{\star}(b,m)=\frac{\beta}{s(b,m)-\beta}\,c_{0}(b,m),$ (41)
and the resulting optimized growth prediction scales as
$\mu_{\mathrm{grow}}^{\star}(B;b,m)\;\approx\;k\,\kappa\big(s(b,m),\beta\big)\,c_{0}(b,m)^{\beta-s(b,m)}\,B^{s(b,m)},$ (42)
where $\kappa(s,\beta)=\beta^{\beta}(s-\beta)^{s-\beta}s^{-s}$. If $s(b,m)\leq\beta$, then (40) is (weakly) increasing in $x$, so the growth approximation favors scale-up over scale-out.
Proposition 12 is a restatement of Theorem 9 in design form. It makes the trade-off explicit: when coordination is expensive (large $c_{0}$), each leaf should be stronger (larger $x^{\star}$); when organization is efficient (large $s-\beta$), each leaf can be weaker, freeing budget to add more leaves.
In practice, given $B$, the recommended number of leaves and depth are
$N^{\star}(B;b,m)\;\approx\;\left\lfloor\frac{B}{x^{\star}(b,m)+c_{0}(b,m)}\right\rfloor,\qquad L^{\star}(B;b,m)\;\approx\;\left\lfloor\log_{b}N^{\star}(B;b,m)\right\rfloor.$ (43)
Depth $L^{\star}$ should not exceed the mixing depth needed to approach saturation (Theorem 6); see Section 6.4.
@@ 6.2 Monotone communication design: why $m^{\star}(B)$ increases with budget
For each $m\in\mathcal{M}_{b}$, define the growth-optimal prediction from Proposition 12,
@@
$t=\log B$$\log\underline{\mu}_{b}(m;B)$upper envelopemm smallmm mediummm large
(a) Envelope over $\log B$
$t=\log B$$m_{b}^{\star}(B)$$m_{1}$$m_{2}$$m_{3}$
(b) Monotone design curve
Figure 10: Why the optimal message length increases with budget in the growth regime (fixed fan-in $b$). After optimizing compute allocation $x$, each candidate message length $m$ induces a line $\log\underline{\mu}_{b}(m;B)=s_{b,m}\log B+q_{b}(m)$ with slope $s_{b,m}$ increasing in $m$. The maximum over $m$ is the upper envelope of these lines (left), implying that an optimal $m_{b}^{\star}(B)$ can be chosen as a nondecreasing, piecewise-constant function of $B$ (right).
@@ 6.4 Putting the diagnostics together
We now combine compute allocation, monotone communication design, and saturation into a compact workflow. The inputs are a budget $B$, a context window $W$, and a small set of estimated effective quantities, in particular $\mu_{0}(x)$, $\rho$, and $\gamma(m)$ (Section 2.7). The output is a model-based prediction and a set of diagnostics (which constraint is binding), rather than a claim that the abstraction optimizes real agent stacks end-to-end.
@@
For each odd fan-in $b\geq 3$ and each feasible $m$ with $bm\leq W$, compute $\alpha_{\rho}(b,m)$ and $s(b,m)$. Discard candidates with $\alpha_{\rho}(b,m)\leq 1$, which collapse in depth (Theorem 4). For exponent-level budgeted synergy in the growth regime, also discard $s(b,m)\leq\beta$. For remaining candidates, set the per-leaf compute to the closed form $x^{\star}(b,m)$ in (41) (Proposition 12) and choose the implied leaf count and depth via (43), with depth interpreted through the mixing-depth guarantee of Theorem 6. For fixed $b$, Algorithm 1 yields the monotone budget-indexed communication curve $m_{b}^{\star}(B)$, so the outer search can be performed efficiently.
@@ 7.2 Synthetic sanity check: recovering the amplification boundary
Synthetic experiments provide a controlled setting where the ground-truth parameters are known and the phase boundary can be probed directly. Our goal here is not to optimize absolute performance, but to verify that the qualitative prediction—amplification versus collapse governed by $\alpha_{\rho}$—is recovered under the generative model analyzed in Section 4.
@@ 8 Discussion and Limitations
8.1 Relationship to prior results
@@
8.2 What the framework explains (and what it does not)
@@
At the same time, the model is not a faithful simulator of real agent stacks. Parameters such as $\rho$ and $\gamma(m)$ are effective summaries that can vary with depth, prompts, tools, and task difficulty; real communication is far richer than a one-bit abstraction; and practical budget accounting can differ across implementations. For these reasons, we view the theory as a coordination diagnostic—predicting which constraint is binding and which direction should help—rather than as a source of precise point predictions. Crucially, the alignment between our derived bottlenecks and independent large-scale observations (Kim et al., 2025) suggests that the effective parameters defined here capture the structural essence of the scaling problem, rendering the theory robust to changes in the underlying base models.
8.3 How stronger protocols fit the framework
@@
8.4 Extensions beyond our scope
@@ 9 Conclusion
Multi-agent systems can exhibit synergy, but under a fixed budget scaling out is not monotone: it can help, saturate, or collapse. We argued that these outcomes become predictable once we make the binding constraints explicit: finite context windows, lossy communication, and shared failures, and summarize an environment by a small set of measurable effective parameters.
In a continuous warm-up we derived closed-form risks for star/chain/tree organizations and exposed correlation and communication floors. For binary success/failure tasks, we proved a sharp amplification–collapse phase transition for deep majority trees and characterized the resulting fixed-point saturation and mixing depth. In the amplifying regime we introduced an organization exponent $s$ governing small-signal growth and showed that budgeted synergy in the growth range occurs exactly when $s>\beta$, yielding closed-form compute allocation and explicit budget thresholds. We also translated these results into simple design diagnostics within the model.
The broader message is that topology is not magic. Hierarchies are useful because context limits can make full centralization infeasible, and they succeed only when communication is sufficiently faithful and shared failures are sufficiently weak. We hope this work helps move agent-system design from ad hoc heuristics toward principled, testable coordination theories.
@@
The authors gratefully acknowledge funding from Canada CIFAR AI Chair Program and the Canada NSERC Discovery Grant (RGPIN-2021-03115).
@@ E.2 Clipped objectives and approximation guarantees
The clipped proxy $\widehat{\mu}_{L}=\min\{\mu^{\star},\mu_{0}\alpha_{\rho}^{L}\}$ in (27) is most naturally interpreted for the regime $\mu_{0}\leq\mu^{\star}$, which includes the small-signal settings used in Section 5.4. In that case, monotonicity and Theorem 4 imply $\mu_{L}\leq\mu^{\star}$ for all $L$, while Lemma 17 implies $\mu_{L}\leq\mu_{0}\alpha_{\rho}^{L}$. Therefore $\mu_{L}\leq\widehat{\mu}_{L}$, justifying the global upper bound claim in Section 4.4. Theorems 5 and 6 provide matching lower bounds in the growth and saturation regimes, respectively.
@@ F.3 Proof of Proposition 8
The comparison between a context-limited star and a supercritical tree in Section 5 is driven by two facts: (i) the star cannot use more than $N_{\max}(m)$ leaves for fixed $m$, hence eventually must spend additional budget on scale-up (which grows like $B^{\beta}$ in the small-signal regime), while (ii) a tree with exponent $s>\beta$ admits a growth regime with $B^{s}$ scaling (Appendix D) before saturating at $\mu^{\star}$ (Appendix E). This yields a provable intermediate budget range where trees can dominate under context constraints, formalized through the sufficient condition $B\geq B_{\mathrm{crit}}$ in Corollary 10.
@@ I.2 Upper-envelope construction and correctness of Algorithm 1
These templates are meant to support qualitative diagnostics: whether context is binding, whether measured correlation suggests limited gains, and whether the calibrated $\widehat{\alpha}_{\rho}$ places a hierarchy in the amplifying or collapsing regime. In that sense, the phase diagrams and design curves from Sections 56 provide concrete hypotheses about which organization should help under matched budgets and which lever (communication fidelity versus shared-failure reduction) is most likely to matter when it does not. Establishing quantitative predictiveness in realistic deployments, where prompts, tools, and dependence structure may drift with depth and task difficulty, is an important direction for future work.

### v2

@@ 4.5.5 A universal upper bound on amplification exponents
Phase-diagram summary and implications.
The preceding results collapse topology choice into a small set of quantitative checks. Under context constraints, deep trees are only viable in the supercritical regime $\alpha_{\rho}(b,m)>1$ (otherwise bias collapses to chance as depth grows). In the growth regime, budgeted synergy against scale-up further requires $s(b,m,\rho)>\beta$, equivalently $\alpha_{\rho}>b^{\beta}$ (Section 4.5.4). Within one-bit majority-vote hierarchies we also have the universal cap $s\leq 1/2$ (Theorem 11), which explains why scale-out is fragile on some tasks: if $\beta\geq 1/2$, exponent-level synergy is impossible without richer communication, heterogeneity that lowers the effective $\rho$, or protocols that change the effective aggregation map. Even when $s>\beta$, gains persist only until saturation at $\mu^{\star}(b,m)$; beyond that point, additional budget is typically better spent improving fidelity (larger $m$, hence larger $\gamma(m)$) or reducing shared failures (smaller $\rho$) than on additional depth. Figure 9 summarizes these regimes, and Section 5 turns the checks into concrete design diagnostics within the model.
@@ 5 Design Diagnostics Within the Model
The results so far characterize when different organizations can help under a fixed budget: a context-limited star saturates (Section 4.5.3), chains compound loss (Section 4.5.2), and trees are viable only in the supercritical regime $\alpha_{\rho}>1$ with small-signal growth governed by an organization exponent $s$ (Section 4). Here we translate these results into design diagnostics within the model: given a budget $B$, a context window $W$, and a small set of effective environment quantities, we ask whether hierarchy is feasible, whether scale-out can outpace scale-up before saturation, and which constraint is binding when it cannot. For LLM-based agent systems, the knobs correspond to token allocations: per-leaf reasoning/compute $x$ and inter-agent message length $m$. The abstraction is deliberately coarse, so the goal is regime prediction and diagnostic guidance, not a faithful end-to-end optimizer for any particular agent stack.
@@
5.1 Compute allocation in the growth regime
Fix a feasible fan-in $b$ and message length $m$ with $bm\leq W$. Let $\alpha_{\rho}(b,m)$ be defined as in (29), and write the associated organization exponent as $s(b,m):=\log\alpha_{\rho}(b,m)/\log b$. In the supercritical regime $\alpha_{\rho}(b,m)>1$ and before saturation, the binary small-signal prediction takes the growth form (41). Theorem 9 yields a closed-form compute allocation $x^{\star}(b,m)$ (eq. (42)) and an optimized growth curve $\mu_{\mathrm{grow}}^{\star}(B;b,m)$ (eq. (43)) whenever $s(b,m)>\beta$. If $s(b,m)\leq\beta$, the growth surrogate favors spending budget to scale up rather than scale out.
Given a budget $B$, we then allocate
$N^{\star}(B;b,m)\approx\left\lfloor\frac{B}{x^{\star}(b,m)+c_{0}(b,m)}\right\rfloor,\qquad L^{\star}(B;b,m)\approx\left\lfloor\log_{b}N^{\star}(B;b,m)\right\rfloor,$ (48)
and cap $L^{\star}$ using the mixing-depth guarantee (Theorem 6) when operating near saturation.
@@ 5.2 Monotone communication design: why $m^{\star}(B)$ increases with budget
For each $m\in\mathcal{M}_{b}$, define the growth-optimal prediction from Theorem 9,
@@
(a) Upper envelope over candidate message lengths.
(b) Resulting $m_{b}^{\star}(B)$ (monotone step function).
Figure 10: Upper-envelope argument behind the monotone message-length design curve (fixed fan-in $b$). For each feasible message length $m$ (numbers shown are tokens in this example), we first optimize the per-leaf compute allocation $x$ in the growth-regime surrogate, yielding a closed-form prediction of the form $\log\underline{\mu}_{b}(m;B)=s_{b,m}\,\log B+q_{b}(m)$, i.e., a line in $t=\log B$ whose slope $s_{b,m}$ increases with $m$ because communication fidelity $\gamma(m)$ improves with longer messages. (a) We plot these candidate lines for a representative subset of feasible $m$ values and highlight their pointwise maximum (thick curve), i.e., the upper envelope. Red markers indicate budgets where two candidates tie (line intersections), and the vertical dashed lines project these switch points onto the $t$-axis. (b) The maximizing message length $m_{b}^{\star}(B)$ is therefore a nondecreasing step function of $B$, jumping exactly at the envelope breakpoints. Algorithm 1 constructs the full envelope efficiently by scanning candidates in increasing $m$ (increasing slope) and maintaining only the active lines and their intersection thresholds.
@@ 5.4 Putting the diagnostics together
We now combine compute allocation, monotone communication design, and saturation into a compact workflow. The inputs are a budget $B$, a context window $W$, and a small set of estimated effective quantities, in particular $\mu_{0}(x)$, $\rho$, and $\gamma(m)$ (Section 2.7). The output is a model-based prediction and a set of diagnostics (which constraint is binding), rather than a claim that the abstraction optimizes real agent stacks end-to-end. Figure 11 summarizes the workflow of design-diagnostics.
Inputs / measured quantities
budget $B$, context $W$, single-agent exponent $\beta$ (or $g(\cdot)$), fidelity curve $\gamma(m)$, shared-failure correlation $\rho$
Feasible design grid
choose odd fan-in $b$ and message length $m$ subject to $bm\leq W$
Amplification diagnostics
compute $\alpha_{\rho}(b,m)$ (29), compute $s(b,m,\rho)$ (34), check $\alpha_{\rho}>1$ and $s>\beta$
Budget split and depth
compute $x^{\star}$ (42), compute $N^{\star},L^{\star}$ (48), cap depth using Theorem 6
Predict and compare
growth prediction $\mu^{\star}_{\mathrm{grow}}$ (43), clip at saturation $\mu^{\star}$ (Sec. 4.4), compare to star and single baselines
Outputs
recommended $(b,m,x^{\star},L^{\star})$, predicted regime: growth vs. saturation diagnostic: which constraint binds
Figure 11: A design-diagnostics pipeline within the model. Measured constraints ($B,W$) and calibrated parameters ($\beta$, $\gamma(m)$, $\rho$) determine which $(b,m)$ are feasible, whether depth amplifies ($\alpha_{\rho}>1$), and whether scale-out can beat scale-up in the growth regime ($s>\beta$). The closed-form allocation $x^{\star}$ and implied $(N^{\star},L^{\star})$ yield a predicted performance that is clipped at the saturation fixed point $\mu^{\star}$ and compared against star and single-agent baselines.
@@
For each odd fan-in $b\geq 3$ and each feasible $m$ with $bm\leq W$, compute $\alpha_{\rho}(b,m)$ and $s(b,m)$. Discard candidates with $\alpha_{\rho}(b,m)\leq 1$, which collapse in depth (Theorem 4). For exponent-level budgeted synergy in the growth regime, also discard $s(b,m)\leq\beta$. For remaining candidates, set the per-leaf compute to the closed form $x^{\star}(b,m)$ in (42) (Theorem 9) and choose the implied leaf count and depth via (48), with depth interpreted through the mixing-depth guarantee of Theorem 6. For fixed $b$, Algorithm 1 yields the monotone budget-indexed communication curve $m_{b}^{\star}(B)$, so the outer search can be performed efficiently.
@@ 6.2 Synthetic sanity check: recovering the amplification boundary
Synthetic experiments provide a controlled setting where the ground-truth parameters are known and the phase boundary can be probed directly. Our goal here is not to optimize absolute performance, but to verify that the qualitative prediction, i.e., amplification versus collapse governed by $\alpha_{\rho}$, is recovered under the generative model analyzed in Section 4.
@@ 7 Discussion and Limitations
While our results provide a rigorous asymptotic foundation for agent scaling, the practical utility of these bounds hinges on the empirical estimation of the shared-failure correlation $\rho$ and the organization exponent $s$. In practice, $\rho$ serves as a ‘homogeneity metric’ for a given model ensemble; a high $\rho$ suggests that the agents are inheriting the same inductive biases or training data flaws, thereby collapsing the effective fan-in $b$. Future work should focus on developing diagnostic probes—short, high-entropy test batteries—to calibrate these parameters for specific LLM families. By doing so, the Phase Transition $\alpha_{\rho}$ can be transformed from a theoretical threshold into a real-time ‘stopping rule’ for determining whether further tree depth will yield diminishing returns or catastrophic signal decay.
Relationship to prior results
@@
What the framework explains (and what it does not)
@@
At the same time, the model is not a faithful simulator of real agent stacks. Parameters such as $\rho$ and $\gamma(m)$ are effective summaries that can vary with depth, prompts, tools, and task difficulty; real communication is far richer than a one-bit abstraction; and practical budget accounting can differ across implementations. For these reasons, we view the theory as a coordination diagnostic, i.e., predicting which constraint is binding and which direction should help, rather than as a source of precise point predictions. Crucially, the alignment between our derived bottlenecks and independent large-scale observations (Kim et al., 2025) suggests that the effective parameters defined here capture the structural essence of the scaling problem, rendering the theory robust to changes in the underlying base models.
How stronger protocols fit the framework
@@
Extensions beyond our scope
@@ 8 Conclusion
Multi-agent systems can exhibit synergy, but under a fixed budget scaling out is not monotone: it can help, saturate, or collapse. We argued that these outcomes become predictable once we make the binding constraints explicit: finite context windows, lossy communication, and shared failures, and summarize an environment by a small set of measurable effective parameters. In a continuous warm-up we derived closed-form risks for star/chain/tree organizations and exposed correlation and communication floors. For binary success/failure tasks, we proved a sharp amplification–collapse phase transition for deep majority trees and characterized the resulting fixed-point saturation and mixing depth. In the amplifying regime we introduced an organization exponent $s$ governing small-signal growth and showed that budgeted synergy in the growth range occurs exactly when $s>\beta$, yielding closed-form compute allocation and explicit budget thresholds. We also translated these results into simple design diagnostics within the model. The broader message is that topology is not magic. Hierarchies are useful because context limits can make full centralization infeasible, and they succeed only when communication is sufficiently faithful and shared failures are sufficiently weak. We hope this work helps move agent-system design from ad hoc heuristics toward principled, testable coordination theories.
@@
The authors gratefully acknowledge funding from Canada CIFAR AI Chair Program and the Canada NSERC Discovery Grant program.
@@ E.2 Clipped objectives and approximation guarantees
The clipped proxy $\widehat{\mu}_{L}=\min\{\mu^{\star},\mu_{0}\alpha_{\rho}^{L}\}$ in (35) is most naturally interpreted for the regime $\mu_{0}\leq\mu^{\star}$, which includes the small-signal settings used in Section 4.5.4. In that case, monotonicity and Theorem 4 imply $\mu_{L}\leq\mu^{\star}$ for all $L$, while Lemma 16 implies $\mu_{L}\leq\mu_{0}\alpha_{\rho}^{L}$. Therefore $\mu_{L}\leq\widehat{\mu}_{L}$, justifying the global upper bound claim in Section 4.4. Theorems 5 and 6 provide matching lower bounds in the growth and saturation regimes, respectively.
@@ F.3 Proof of Proposition 8
The comparison between a context-limited star and a supercritical tree in Section 4.5 is driven by two facts: (i) the star cannot use more than $N_{\max}(m)$ leaves for fixed $m$, hence eventually must spend additional budget on scale-up (which grows like $B^{\beta}$ in the small-signal regime), while (ii) a tree with exponent $s>\beta$ admits a growth regime with $B^{s}$ scaling (Appendix D) before saturating at $\mu^{\star}$ (Appendix E). This yields a provable intermediate budget range where trees can dominate under context constraints, formalized through the sufficient condition $B\geq B_{\mathrm{crit}}$ in Corollary 10.
@@ I.2 Upper-envelope construction and correctness of Algorithm 1
These templates are meant to support qualitative diagnostics: whether context is binding, whether measured correlation suggests limited gains, and whether the calibrated $\widehat{\alpha}_{\rho}$ places a hierarchy in the amplifying or collapsing regime. In that sense, the phase diagrams and design curves from Sections 4.55 provide concrete hypotheses about which organization should help under matched budgets and which lever (communication fidelity versus shared-failure reduction) is most likely to matter when it does not. Establishing quantitative predictiveness in realistic deployments, where prompts, tools, and dependence structure may drift with depth and task difficulty, is an important direction for future work.